# Learning the Hamiltonian Matrix of Large Atomic Systems

### v1

@@ Appendix I Additional Tests
I.1 Using two HfO2 structures for training
While training on a single structure is sufficient to achieve the results in Fig. 5, increasing the quantity of training data leads to improved prediction accuracy. To demonstrate this, we train using augmented partitioning on two structures, including structure 2 from Table 1 as well as a newly generated structure ”0” of similar size (3000 atoms). Training was performed with 18 slices of 3 Å  thickness taken from both structures, validated on structure 1, and tested on structure 3. The results are compared against the previous model trained on one structure in Table 9, showing that the prediction accuracy can be improved from 5.87 $meV$ to 3.55 $meV$.
Structures Epochs $\mathbf{\epsilon_{n}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{e}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{tot}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{tot}}\left[meV\right]$
Training Testing
0 3 15,675 2.29 0.20 0.22 5.87
0, 2 3 48,356 2.22 0.12 0.13 3.55
Table 9: Comparison of models trained on one HfO2 structure (structure 2) vs two HfO2 structures (0 and 2). Both are tested on a fully unseen structure (3), showing improved accuracy from 5.87 to 3.55 $meV$.
I.2 Sub-stoichiometric hafnium oxide
Here, we provide a more detailed study on the prediction of substoichiometric HfO2, where the train/test structures contain different vacancy fractions. 18 slices (tslice = 3 Å) were used in all cases. Note that for this experiment, the maximum number of epochs was fixed at 20,000. The results are summarized in Table 10. The $\epsilon_{n}$ and $\epsilon_{e}$ values across different experiments lie within a small range (2.50-2.90 $mE_{h}$ and 0.16-10.18 $mE_{h}$ respectively), showing that the network generalizes well to structures of different vacancies, regardless of which vacancy configuration it was trained on. To demonstrate that the augmented partitioning approach similarly does not affect accuracy for sub-stoichiometric HfOx, we also perform full graph training using structure 2 with 15% vacancies, and compare with the augmented partitioning approach in Table 12. The minimal difference in $\epsilon_{n}$ and $\epsilon_{e}$ values between full and partitioned approaches indicates that both approaches generalize equally well to different stoichiometry. These values are also close to that of stoichiometric HfO2 in Table 3.
@@ I.2 Sub-stoichiometric hafnium oxide
I.3 Partitioning tests for other materials
For each material, we train the model on partitions of different thicknesses containing the same number of atoms. The results are summarized in Table 11.
@@ I.3 Partitioning tests for other materials
a-GST 30 1 12 710 138532 1.10 0.09 0.09 2.53
a-GST 5 6 12 710 48356 0.97 0.10 0.10 2.77
a-PtGe 50 1 8 1633 182372 0.81 0.08 0.09 2.37
a-PtGe 5 10 8 1633 84416 0.78 0.08 0.09 2.39
Table 11: Summary of model performance when trained on amorphous GST and PtGe structure with slices of different thicknesses. The structures used for training, validation, and testing are displayed in Table 1.
The same test was also repeated on substoichiometric HfO1.7 and tested on structures with different stoichiometry (shown in Table 12. It can be seen that augmented partitioning preserves the accuracy of prediction for all material cases studied. For HfOx, it also preserves the model’s ability to generalize to structures of different stoichiometry.
@@
partitioned 1.9 2.94 0.16
partitioned 1.8 2.52 0.16
partitioned 1.7 2.52 0.16
full 1.9 2.96 0.19
full 1.8 2.67 0.18
full 1.7 2.64 0.17

### v2

@@ Appendix I Additional Tests
I.1 Sub-stoichiometric hafnium oxide
Here, we provide a more detailed study on the prediction of substoichiometric HfO2, where the train/test structures contain different vacancy fractions. 18 slices (tslice = 3 Å) were used in all cases. The results are summarized in Table 10. The $\epsilon_{n}$ and $\epsilon_{e}$ values across different experiments lie within a small range (2.24-2.73 $mE_{h}$ and 0.13-10.18 $mE_{h}$ respectively), showing that the network generalizes well to structures of different vacancies, regardless of which vacancy configuration it was trained on. To demonstrate that the augmented partitioning approach similarly does not affect accuracy for sub-stoichiometric HfOx, we also perform full graph training using structure 2 with 15% vacancies, and compare with the augmented partitioning approach in Table 12. The comparison of $\epsilon_{n}$ and $\epsilon_{e}$ values between full and partitioned approaches indicates that both approaches generalize well to different stoichiometry. These values are also close to that of stoichiometric HfO2 in Table 3.
@@ I.1 Sub-stoichiometric hafnium oxide
I.2 Partitioning tests for other materials
For each material, we train the model on partitions of different thicknesses containing the same number of atoms. The results are summarized in Table 11. The same test was also repeated on substoichiometric HfO1.7 and tested on structures with different stoichiometry (shown in Table 12. It can be seen that augmented partitioning preserves the accuracy of prediction for all material cases studied. For HfOx, it also preserves the model’s ability to generalize to structures of different stoichiometry.
@@ I.2 Partitioning tests for other materials
a-GST 30 1 12 710 138532 1.15 0.08 0.08 2.28
a-GST 5 6 12 710 48356 0.97 0.08 0.09 2.40
a-PtGe 50 1 8 1633 182372 0.77 0.08 0.08 2.26
a-PtGe 5 10 8 1633 84416 0.77 0.08 0.09 2.49
Table 11: Summary of model performance when trained on amorphous GST and PtGe structure with slices of different thicknesses. The structures used for training, validation, and testing are displayed in Table 1. $N_{n}$ and $N_{t}$ refers to the total number of labeled nodes and edges in the training set respectively.
@@
partitioned 1.9 2.32 0.14
partitioned 1.8 2.37 0.15
partitioned 1.7 2.51 0.18
full 1.9 3.02 0.12
full 1.8 2.54 0.11
full 1.7 2.52 0.11
@@
I.3 Small Molecule Benchmarks
Dataset Model Train Validate Test Batch size $\mathbf{\epsilon_{tot}}\left[\mu E_{h}\right]$
water QHNet 500 500 3900 10 10.79
This work 500 500 3900 10 5.60
malondialdehyde QHNet 25,000 500 1,478 5 21.52
This work 25,000 50 1,478 5 16.60
uracil QHNet 25,000 500 4,500 5 20.12
This work 25,000 50 4,500 5 13.80
Table 13: Comparison with QHNet on MD17 benchmark dataset containing water, malondialdehyde and uracil
We test our backbone architecture setup on part of the MD17 benchmark (Schütt et al., 2019), which consists of small-molecules, each with 3 (water), 9 (malondialdehyde) or 12 (uracil) atoms. The hyperparameters used are listed in Table. 14. The number of training steps for each experiment is fixed at 200,000 to match that of QHNet in (Yu et al., 2023b) . The results in Table 13 show that the performance of our setup is comparable to that of the baseline for all measured cases.
Hyper-parameters MD 17 dataset
Optimizer Adam
Precision double (f64)
Scheduler ReduceLROnPlateau
Criterion RMSE + MAE
Initial learning rate $1\times 10^{-3}$
Minimum learning rate $1\times 10^{-10}$
Number of MP Layers 2
Decay patience $t_{patience}$ $50$
Decay factor $0.5$
Threshold $1\times 10^{-5}$
Maximum degree $L_{max}$ $4$
Maximum order $M_{max}$ $4$
Embedding size $128$
Number of attention heads $N_{h}$ $2$
Feedforward Network Dimension $64$
Table 14: Hyper-parameters used for MD17 dataset.